# Neural Force Field: Learning Generalized Physical Representation from a Few Examples

### v1

@@ B.1 Hyperparameters
In I-PHYRE, we train the models using a single NVIDIA A100 Tensor Core GPU. The force field predictor in NFF is based on a DeepONet architecture, consisting of trunk and branch networks, each a 3-layer MLP with a hidden size of 256. The ODE solver uses the Euler method with a step size of 0.005. The 150-step trajectories are divided into 6-step segments, with 6 trajectories per batch. The learning rate starts at 5e-4 and gradually decays to 1e-5 following a cosine annealing schedule. Training occurs over 3000 epochs with a weight decay of 1e-5 for regularization.
@@ B.2 Evaluation metrics
FPE$=\left|\hat{z}_{\text{final}}-z_{\text{final}}\right|.$ (A2)
@@
R$=\frac{\sum_{t=1}^{n}(\hat{z}_{t}-\bar{\hat{z}})(z_{t}-\bar{z})}{\sqrt{\sum_{t=1}^{n}(\hat{z}_{t}-\bar{\hat{z}})^{2}\sum_{t=1}^{n}(z_{t}-\bar{z})^{2}}}.$ (A4)
Appendix C Overfitting
Figure A1: SlotFormer overfitting on physical data. While the performance improves in the within-scenario setting as training progresses, the cross-scenario setting experiences overfitting, resulting in increasing RMSE during training.
We demonstrate that methods without interaction modeling are prone to overfitting on in-distribution physical data. While they perform well within the same scenario, they face significant challenges in cross-scenario settings. To illustrate this, we train a SlotFormer on N-body training data and validate it on a separate 4-body cross-scenario dataset. As shown in Figure A1, the increase in overfitting leads to a decline in generalization performance.
Appendix D Ablation studies on integration methods
In Table A1, we present additional ablation studies comparing various integration methods. Specifically, we evaluate the performance of different integration orders on the N-body task, including Euler, Midpoint, Heun3, RK4, and adaptive methods. Computational complexity is measured by the average number of integration steps. The results indicate that Euler integration provides the best cross-scenario generalization, while Heun3 excels in within-scenario generalization. Despite this, Euler integration has lower computational complexity than the higher-order methods. These findings suggest that Euler integration is sufficient for our physical reasoning tasks.
Table A1: Results of different kinds of integration methods. The computational complexity of the model is proportional to the number of average integration steps (relative to Euler $35e-$ 3). Among non-adaptive integration methods, the Euler method achieves the best generalization performance with relatively low computational complexity, whereas adaptive methods at different tolerance levels do not exhibit consistent performance improvements.
@@ Appendix D Ablation studies on integration methods
Adaptive ($31e-$ 3) $0.20$ $0.0725\pm 0.0084$ $0.1537\pm 0.0069$
Adaptive ($41e-$ 4) $0.28$ $0.0582\pm 0.0035$ $0.1797\pm 0.0125$
Adaptive ($51e-$ 5) $0.36$ $0.0783\pm 0.0056$ $0.1539\pm 0.0056$
Adaptive ($61e-$ 6) $0.56$ $0.0807\pm 0.0053$ $0.1668\pm 0.0067$
Adaptive ($71e-$ 7) $1.07$ $0.0811\pm 0.0053$ $0.1928\pm 0.0077$
@@
Euler ($35e-$ 3) $1.00$ $0.0785\pm 0.0069$ $0.1522\pm 0.0066$
Midpoint ($35e-$ 3) $2.00$ $0.0671\pm 0.0053$ $0.1762\pm 0.0071$
Heun3 ($35e-$ 3) $3.00$ $0.0606\pm 0.0041$ $0.1550\pm 0.0067$
RK4 ($35e-$ 3) $4.00$ $0.0682\pm 0.0048$ $0.1626\pm 0.0072$
Appendix E Detailed planning results
We present the detailed planning results of different models across trials in Table A2.
@@ Appendix E Detailed planning results
We test the performance of GPT-4o for both within-scenario planning and cross-scenario planning. Detailed results are presented in Table A3. Within-scenario games contains the same 10 games used in the prompt in Round 1. So the language model can memorize corresponding solutions provided in the prompt to achieve good performance. Cross-scenario games contains the other 30 game settings, on which the performance of GPT-4o is better than other baselines (Random, IN, SlotFormer) in Table A2. The improvement of performance through refinement is witnessed in both within-scenario and cross-scenario games.
@@
NFF $\times$ $0.51\pm 0.08$ $0.55\pm 0.08$ $0.58\pm 0.08$ $0.60\pm 0.08$ $0.62\pm 0.08$
NFF $\checkmark$ $0.51\pm 0.08$ $0.62\pm 0.08$ $0.69\pm 0.07$ $0.74\pm 0.06$ $0.77\pm 0.06$
Human $\times$ $0.52\pm 0.06$ $0.66\pm 0.06$ $0.73\pm 0.06$ $0.77\pm 0.06$ $0.80\pm 0.05$
@@
The prompt used in Round 1.
Appendix F More visualization
We present additional visualization of learned forces along the trajectories in I-PHYRE in Figures A2 and A3. We also present more prediction results on N-body in Figure A4.
@@ Appendix F More visualization
(a) Ground truth trajectory
(b) Our NFF prediction
(c) SlotFormer (Wu et al., 2023)
(d) IN (Battaglia et al., 2016)
Figure A4: Additional prediction results on N-body.

### v4

@@ B.1 Hyperparameters
In I-PHYRE, we train the models using a single NVIDIA A100 Tensor Core 80GB GPU. The force field predictor in NFF is based on a DeepONet architecture, consisting of trunk and branch networks, each a 3-layer MLP with a hidden size of 256. The ODE solver uses the Euler method with a step size of 0.005. The 150-step trajectories are divided into 6-step segments, with 6 trajectories per batch. The learning rate starts at 5e-4 and gradually decays to 1e-5 following a cosine annealing schedule. Training occurs over 3000 epochs with a weight decay of 1e-5 for regularization.
@@
For training the vision-based NFF in PHYRE, we perform feature extraction from $256\times 256$ object masks using four sparse convolutional layers with channel sizes of 16, 32, 32, and 32. The model architecture includes a DeepONet, composed of a trunk network and a branch network, each with 3 layers and 128 hidden units. DeepONet takes as input a combination of mask, positional, and velocity features, and outputs the corresponding 2D forces and torques acting on each object. Then, we perform a transform and rotation on the object masks to obtain the frames in the next states. The ODE solver uses the Euler method with a step size of 0.025. We adopt a cosine annealing schedule to adjust the learning rate from 3e-4 to 5e-5, with a batch size of 8. Training is conducted for 500 epochs to ensure convergence, with the 18-step trajectories divided into 2-step segments during training.
@@ B.2 Evaluation metrics
[…]
@@
[…]
Appendix C Ablation studies on integration methods
In Tab.˜A1, we present additional ablation studies comparing various integration methods. Specifically, we evaluate the performance of different integration orders on the N-body task, including Euler, Midpoint, Heun3, RK4, and adaptive methods. Computational complexity is measured by the average number of integration steps. The results indicate that Heun3 excels in within-scenario and cross-scenario generalization. Despite this, Euler integration has lower computational complexity than the higher-order methods, while maintaining better training stability compared to adaptive methods. These findings suggest that Euler integration is sufficient for our physical reasoning tasks.
Table A1: Results of different kinds of integration methods. The computational complexity of the model is proportional to the number of average integration steps (relative to Euler 5e35e-3). Among non-adaptive methods, with the increase of computational complexity, the Heun3 method achieves the best performance. For uniformity across tasks and stability during training, we simply selected the Euler method for integration.
@@ Appendix C Ablation studies on integration methods
Adaptive (1e41e-4) $0.28$ $0.091\pm 0.008$ $1.100\pm 0.021$
Adaptive (1e51e-5) $0.36$ $0.093\pm 0.008$ $1.063\pm 0.020$
Adaptive (1e61e-6) $0.56$ $0.107\pm 0.008$ $1.082\pm 0.023$
Adaptive (1e71e-7) $1.07$ $0.119\pm 0.009$ $1.125\pm 0.026$
@@
Euler (5e35e-3) $1.00$ $0.139\pm 0.013$ $1.090\pm 0.023$
Midpoint (5e35e-3) $2.00$ $0.091\pm 0.009$ $1.040\pm 0.023$
Heun3 (5e35e-3) $3.00$ $0.091\pm 0.008$ $1.022\pm 0.022$
RK4 (5e35e-3) $4.00$ $0.103\pm 0.009$ $1.084\pm 0.029$
Appendix D Comparison of baselines under noise
We evaluate the robustness of our model under noisy conditions by introducing noise at varying levels to the N-body simulated trajectories during training. As shown in Tab.˜A2, NFF demonstrates consistently strong performance across both within-scenario and cross-scenario settings. In contrast, IN tends to diverge significantly from the ground truth, particularly in the cross-scenario case, where it exhibits large errors.
Additionally, to simulate imperfect perception in I-PHYRE, we extract object states from raw RGB images using OpenCV. The results in Tab.˜A3 indicate that NFF continues to perform better than other models in zero-shot planning scenarios.
Table A2: NFF demonstrates stronger robustness than other models under imperfect perception. For N-body, we introduced random noise of 1% to 3% to the training data to mimic imperfect simulation.
Model Within Cross
w/o noise 1% noise 3% noise w/o noise 1% noise 3% noise
IN 0.31 $\pm$ 0.02 0.33 $\pm$ 0.03 0.50 $\pm$ 0.03 4.73 $\pm$ 0.11 2.71 $\pm$ 0.06 8.63 $\pm$ 0.78
SlotFormer 0.26 $\pm$ 0.02 0.26 $\pm$ 0.02 0.37 $\pm$ 0.03 3.53 $\pm$ 0.13 3.49 $\pm$ 0.15 2.29 $\pm$ 0.04
NFF 0.14 $\pm$ 0.01 0.12 $\pm$ 0.01 0.13 $\pm$ 0.01 1.09 $\pm$ 0.02 1.32 $\pm$ 0.04 1.53 $\pm$ 0.03
Table A3: Probability of success in I-PHYRE under noisy perception. We use OpenCV to extract states from raw RGB images and then evaluate the model’s zero-shot success rate on cross-scenario games.
Model Ground-truth states RGB images
Random 0.24 $\pm$ 0.04 0.24 $\pm$ 0.04
IN 0.25 $\pm$ 0.06 0.25 $\pm$ 0.06
SlotFormer 0.30 $\pm$ 0.07 0.23 $\pm$ 0.06
NFF 0.51 $\pm$ 0.08 0.41 $\pm$ 0.08
Appendix E Planning time on N-body
We demonstrate how backward integration accelerates planning in N-body tasks. As shown in Tab.˜A4, backward planning is up to 10 times faster than traditional optimization-based planning methods.
Table A4: NFF enables fast planning through backward integration. NFF performs planning approximately 10 times faster than optimization-based methods in the N-body task. The value represents the time in seconds required to plan 200 scenarios in parallel, with each experiment conducted three times.
Model Type Within Cross
IN Optimization 13.886 $\pm$ 0.225 11.693 $\pm$ 0.153
SlotFormer Optimization 22.098 $\pm$ 0.156 22.489 $\pm$ 0.068
NFF Backward integration 1.717 $\pm$ 0.023 1.716 $\pm$ 0.013
Appendix F Detailed planning results
We present the detailed planning results of different models across trials in Tab.˜A5.
@@ Appendix F Detailed planning results
We test the performance of GPT-4o for both within-scenario planning and cross-scenario planning. Detailed results are presented in Tab.˜A6. Within-scenario games contains the same 10 games used in the prompt in Round 1. So the language model can memorize corresponding solutions provided in the prompt to achieve good performance. Cross-scenario games contains the other 30 game settings, on which the performance of GPT-4o is better than other baselines (Random, IN, SlotFormer) in Tab.˜A5. The improvement of performance through refinement is witnessed in both within-scenario and cross-scenario games.
@@
NFF $\times$ $1\pm 0.08 0.55±0$$5\pm 0.08$ $0.58\pm 0.08$ $0.60\pm 0.08$ $0.62\pm 0.08$
NFF $\checkmark$ $0.51\pm 0.08$ $0.62\pm 0.08$ $0.69\pm 0.07$ $4\pm 0.06 0.77±0$$7\pm 0.06$
Human $\times$ $0.52\pm 0.06$ $0.66\pm 0.06$ $3\pm 0.06 0.77±0$$7\pm 0.06$ $0.80\pm 0.05$
@@
The prompt used in GPT-4o.
Appendix G More visualization
We provide additional visualizations of the learned forces along trajectories in I-PHYRE, as shown in Fig.˜A1, Fig.˜A2, and Fig.˜A3. Further prediction results on the N-body dataset are presented in Fig.˜A4. Additionally, examples of failure cases in NFF predictions are shown in Fig.˜A6. We also visualize the failure encoding of slot-attention in SlotFormer in Fig.˜A8.
(a) Support interaction
(b) Platform orientation
(c) Impact dynamics
(d) Spring mechanics
Figure A1: Force response analysis under controlled state variations. Each plot shows the predicted force components (Fx in blue, Fy in orange) from our trained NFF model, measured in normalized units against different state parameters. The subplots systematically analyze: (a) support forces as a function of ball height, showing characteristic contact response when the ball meets the platform, (b) force decomposition as the platform rotates from horizontal ($0^{\degree}$) to vertical ($90^{\degree}$), (c) impact forces scaling with the ball’s downward velocity, and (d) spring forces varying with horizontal displacement from the equilibrium position. Each plot varies one parameter while keeping all other scene variables constant, demonstrating NFF’s learned physical principles, including contact mechanics, geometric reasoning, impact dynamics, and harmonic motion.
@@ Appendix G More visualization
(a) Ground truth
(b) Our NFF
(c) SlotFormer
(d) IN
Figure A4: Additional prediction results on N-body. The figure displays predictions from simple N-body dynamics to complex dynamics with more bodies and extended prediction horizons.
True
NFF
(a) (b) (c)
Figure A6: Examples of prediction failures in PHYRE cross-scenario setting. (a) Minor prediction inaccuracies accumulate, causing the green ball to fall out of the cup unexpectedly. (b) Extremely tiny red balls introduce unforeseen interactions, potentially resulting in significant dynamic changes, such as cup rotation. (c) Complex interactions between the ball and the seesaw can hinder accurate force prediction.
True
Slot-attention
(a) (b) (c)
Figure A8: Examples of failures in Slotformer’s pretrained slot-attention module in cross-scenario setting. The pretrained slot-attention encoder in slotformer, responsible for encoding visual features, exhibits poor generalization to cross-scenario scenes, leading to an inability to capture the precise geometric information crucial for physical reasoning.